OpenStudy (anonymous):

There were 44,000 people at a ball game in Los Angeles. The day's receipts were $ 369,000. How many people paid $ 13.00 for reserved seats and how many paid $ 6.00 for general admission?

OpenStudy (anonymous):

I saw this before

OpenStudy (anonymous):

like the exact same question was asked last week , probably just different numbers lol

OpenStudy (dumbcow):

x,y is num of people in reserved seats and general seats respectively in thousands x+y = 44 13x + 6y = 369 multiply top equation by -6 -6x -6y = -264 13x +6y = 369 _________________ 7x +0 = 105 x = 15 y = 44-15 = 29

OpenStudy (anonymous):

so 29,000 people paid 6.00 and 15,000 people paid 13.00?

OpenStudy (dumbcow):

correct :)

OpenStudy (anonymous):

thank you very much....word problems always mess me up.